# Reflection-Reinforced Self-Training for Language Agents

### v1

@@ Abstract
Self-training can potentially improve the performance of language agents without relying on demonstrations from humans or stronger models. The general process involves generating samples from a model, evaluating their quality, and updating the model by training on high-quality samples. However, self-training can face limitations because achieving good performance requires a good amount of high-quality samples, yet relying solely on model sampling for obtaining such samples can be inefficient. In addition, these methods often disregard low-quality samples, failing to leverage them effectively. To address these limitations, we present Reflection-Reinforced Self-Training (Re-ReST), which leverages a reflection model to refine low-quality samples and subsequently uses these improved samples to augment self-training. The reflection model takes both the model output and feedback from an external environment (e.g., unit test results in code generation) as inputs and produces improved samples as outputs. By employing this technique, we effectively enhance the quality of inferior samples, and enrich the self-training dataset with higher-quality samples efficiently. We perform extensive experiments on open-source language agents across tasks, including multi-hop question answering, sequential decision-making, code generation, visual question answering, and text-to-image generation. Results demonstrate improvements over self-training baselines across settings. Moreover, ablation studies confirm the reflection model’s efficiency in generating quality self-training samples and its compatibility with self-consistency decoding. Our code is released at https://github.com/PlusLabNLP/Re-ReST.
@@ 1 Introduction
Large language models (LLMs) (Kenton & Toutanova, 2019; Touvron et al., 2023; Achiam et al., 2023) have demonstrated potential in interacting with external environments and addressing practical interactive tasks, resulting in a new class of language agents (Nakano et al., 2021; Yao et al., 2022). Many previous methods for training these agents rely on data generated by stronger models (e.g., GPT-4) (Chen et al., 2023; Yin et al., 2024). However, the reliance on stronger models, which are not always available, limits the agents’ performance. Among the potential techniques to improve their performance (Ouyang et al., 2022; Wang et al., 2023b; Li et al., 2024; Chen et al., 2024), self-training holds promise for improving model performance on unlabeled data without being constrained by the availability and quality of demonstrations from humans or stronger models.
Self-training generally involves refining the model by generating samples, assessing their quality through rewards, and updating the model by training on high-quality samples. Among this line of research (He et al., 2019; Huang et al., 2023; Dong et al., 2023), Reinforced Self-Training (ReST) (Gulcehre et al., 2023) ranks and filters the generated samples with a scoring function, and the selected samples are added to the training data. Within a similar paradigm, RL from AI feedback (RLAIF) Bai et al. (2022); Lee et al. (2023) and self-rewarding (Yuan et al., 2024) uses the LLM itself as a judge to score each generated sample, and trains the model with reinforcement learning or direct preference optimization (DPO) (Rafailov et al., 2024) on the scored samples; STaR (Zelikman et al., 2022) specifically targets the reasoning ability of LLMs and asks an LLM to generate rationales for self-training, enabling the LLM to learn and improve from its own generated reasoning.
@@
However, these self-training methods have several limitations. One significant challenge lies in the acquisition of high-quality samples necessary for achieving good performance. To illustrate, self-training requires a good amount of high-quality samples, but relying solely on model-generated samples can be inefficient, especially for challenging tasks and scenarios where it is hard to find good samples only through sampling. Moreover, the common practice of discarding low-quality samples neglects their potential for improvement and effective utilization, thus limiting the overall efficacy of self-training methods.
Using LLMs to reflect upon their generated samples and perform corrections automatically can be a promising solution to address the issues. Several works (Madaan et al., 2023; Shinn et al., 2023; Pan et al., 2023) have showcased the self-improvement ability of LLMs, particularly emphasizing their ability to refine performance during inference without further model training. However, as pointed out by Huang et al. (2024), high-quality external feedback is necessary for LLMs to correctly improve their generations, which is often unavailable during the test time. In addition, performing self-reflection can introduce computational overhead during inference.
To address these issues, we propose Reflection-Reinforced Self-Training (Re-ReST), which aims to improve the self-training algorithm with a reflection model. Specifically, as in Madaan et al. (2023); Shinn et al. (2023), we employ a reflection model to correct low-quality samples. The reflection model is given the model samples as well as feedback from environments (e.g. if an execution succeeded and unit test results), which is often available during training and thus making it possible for the reflector to refine the model samples given the additional information. This way, the reflector can turn low-quality samples into high-quality ones, enriching the training data for bootstrapping. After the training phase, only the agent model is utilized for inference, ensuring zero additional computational burden during the test time.
We perform extensive experiments with open-source LLMs on a wide range of tasks, including multi-hop question answering, sequential decision-making, code generation, visual question answering, and text-to-image generation. Results show consistent improvements over vanilla self-training across tasks, demonstrating the effectiveness of our method. In addition, ablation studies and analyses confirm that using a reflector can obtain high-quality samples more efficiently than only relying on model sampling, and our training method is compatible with self-consistency decoding.
@@ 2 Method: Re-ReST
Preliminary: Self-Training.
Formally, given an unlabeled dataset , classic self-training algorithms start from using a base model to generate a pseudo label for each instance . Afterward, a subset of is selected based on a certain scoring function and is finetuned on this selected subset. This process can be repeated until model convergence or maximum iterations are reached.
@@
We present Re-ReST, which aims to improve the pseudo-label generation process of self-training. In the self-training paradigm, if a generated pseudo label is of low quality, it will be filtered because including it is not helpful or even harmful for self-training. To address this issue, as shown in Figure 1, we propose to improve low-quality samples with a reflection model given external feedback. Then, we enrich the self-training data by incorporating the corrected generations.
@@ 2.1 Components
The language agent is built upon a large language model (LLM) that is trained or prompted to generate the text and actions given a task. Formally, given an instance , the agent generates its output containing its actions. The agent can optionally first generate its reasoning traces before outputting its actions, which has been demonstrated to improve the model performance and interpretability (Yao et al., 2022).
@@
The reflection model is also instantiated as an LLM, the goal of which is to improve the language agent’s generations given external feedback. We assume that during training, an external environment can evaluate a generated sample and provide feedback to the agent. The feedback can be a binary success status and/or error information. For example, in code generation tasks, the environment can execute the model-generated code on unit tests, providing information on whether the code has syntax errors and whether it can pass the unit tests. Having access to such an environment is important in our setting, as it has been shown that an LLM cannot perform self-correction without high-quality external feedback (Huang et al., 2024). The reflection model generates a corrected sample given the task information , the agent generation , and the environmental feedback . It can optionally first state its reasoning process (e.g., which specific actions could be corrected) before generating the corrected answer.) The use of the reflection model can improve self-training by finding good solutions efficiently because of the additional information provided (i.e., the agent’s previous trial and the environmental feedback.) Note that we do not share the model parameters between the agent and reflector.
@@ 2.2 Data Generation
We then describe how we generate self-training data for the language agent . The data generation process involves two steps, including the initial generation step with the language agent itself and the reflection step with the reflector, and we will obtain the agent-generated dataset and reflector-generated dataset from the two steps respectively.
@@
As in the standard self-training setup, given an instance , we sample generations from the current language agent model . Then, the environment scores the generation and provides feedback . If the score exceeds a threshold, we add the instance to to the training data . In practice, we observe that setting can achieve a good balance between efficiency and effectiveness as we will show in the experiment section.
@@
The initial generation step only relies on the agent model itself to generate data. For a sampled generation , if the score does not pass the threshold, we will feed it to the reflection model for refinement. The reflector takes as inputs the task information , the agent’s prior generation , and the environmental feedback , and then generates the corrected sample The corrected sample will also be evaluated by the environment and we will add it to the reflector-generated training dataset if its score exceeds the threshold. It is important to note that while the reflection procedure can be iteratively applied multiple times as per Shinn et al. (2023), in this study, we limit this process to a single iteration for the sake of efficiency. This means that each generated sample is allowed a maximum of one refined counterpart .
@@ 2.3 Model Training and Inference
We propose to first train the reflector model parameterized by and then use the trained reflector to generate the reflection data . Afterward, we combine and the agent’s self-generated data to train the agent model parameterized by .
@@
While base LLMs can perform self-reflection or self-correction without any finetuning (Shinn et al., 2023), we propose to further improve its reflection ability with the self-generated data. First, from the initial generation step, we obtain multiple generations from the agent model . For each correct generation and incorrect generation with its environmental feedback in , we will add the instance to the agent-generated dataset for reflector training. In addition, the reflector generates its self-training dataset in a zero-shot manner similar to the agent initial generation step. Combining the two generated datasets, we train the reflector on with the standard maximum log-likelihood objective first before generating the training data for the language agent:
@@
After we have the base language agent to generate the self-training data and the improved reflector to generate the reflector-generated data , we can train the language agent jointly on with maximum log-likelihood:
Besides the maximum log-likelihood objective, because the reflection training and data generation process involves the use of preference pairs, it is natural to use preference optimization objectives such as DPO (Rafailov et al., 2024) for training. We will discuss the use of preference optimization objectives in the experiment section.
@@
During inference, it is often hard to have access to high-quality environmental feedback, which can make inference-time self-reflection algorithms fail (Huang et al., 2024). Therefore, we only have the language agent directly outperform generations without using the reflection model during inference. This way, we do not introduce any computational overhead at the test time. A possible approach to integrate the reflector into inference is to first train a scorer that scores the agent generation, and then perform self-correction if the score is lower than a threshold, which we leave as a future direction.
@@ 3 Experiments
We experiment with multi-hop reasoning, sequential decision-making, code generation, visual question answering, and text-to-image generation to demonstrate the general applicability of Re-ReST. We present the experimental settings and results for each task. In all our experiments, we advocate for the use of open-source models and aim to avoid black-box, closed-source commercial models whenever possible.
@@ 3.1 Multi-Hop Reasoning
We use the HotpotQA dataset (Yang et al., 2018), which is a well-established question-answering dataset featuring multi-hop reasoning and knowledge retrieval. It is constructed based on Wikipedia and an agent needs to retrieve and reason over multiple supporting documents to answer a question. We sample 5,000 training instances randomly for finetuning models and 500 instances from the development set for evaluation as in Chen et al. (2023).
@@
We use 2-shot prompting for zero-shot agent and reflector data generation as in Shinn et al. (2023). For each training instance, the agent model samples 3 generations. The generation is evaluated with the exact match metric (i.e., if the generated answer is exactly the same as the ground-truth answer). The retrieval and evaluation results are given to the reflector as the environmental feedback for self-correction. We use Low-Rank Adaptation (LoRA) (Hu et al., 2022) for training the language models for efficiency. The agent and reflector models are trained for 3 epochs with a learning rate of 3e-4.
| Model | QA Tool | #Train Data | EM (%) |
@@
| Llama-2-13B Agents | |||
| Zero-Shot | WikipediaAPI | - | 20.0 |
@@
| Llama-3-8B Agents | |||
| Zero-Shot | WikipediaAPI | - | 30.0 |
@@
| Llama-2-13B Agents w/ GPT-4-Generated Data | |||
| FireAct (Chen et al., 2023) | SerpAPI | 0/2k | 34.4 |
| LUMOS (Yin et al., 2024) | GPT-4 | 0/20k | 36.3 |
| Self-Training | WikipediaAPI | 2.5k/0.5k | 34.2 |
| Re-ReST | WikipediaAPI | 3k/0.5k | 35.8 |
@@
We list the main results in Table 1. As shown in the table, self-training can significantly improve the model performance from an EM score of 20.0% to 27.6% for Llama-2 and from 30.0% to 34.4% for Llama-3. However, only 37.1% and 48.3% of the training instances are correctly solved by the agent model and are used for self-training respectively. By integrating our reflector model into the process, the agent can solve more training instances and thus have more data for training the agent model, increasing the EM scores significantly. In addition to our implemented models, following previous work (FireAct (Chen et al., 2023) and LUMOS Yin et al. (2024)) that use GPT-3.5/4 for data generation and model fine-tuning, we employ GPT-4 to generate 0.5k instances and first train the agents with the GPT-4 generated data before self-training. Results demonstrate comparable or better performance of our model than these methods, even though both of them use strong knowledge retrieval models (i.e., SerpAPI222https://serpapi.com/ for FireAct and GPT-4 for LUMOS), which are costly and non-scalable. By contrast, we use the free Wikipedia API.

### v3

# Re-ReST: Reflection-Reinforced Self-Training for Language Agents

@@ Abstract
Finetuning language agents with reasoning-action trajectories is effective, but obtaining these trajectories from human annotations or stronger models is costly and sometimes impractical. In this paper, we investigate the use of self-training in language agents, which can generate supervision from the agent itself, offering a promising alternative without relying on human or stronger model demonstrations. Self-training, however, requires high-quality model-generated samples, which are hard to obtain for challenging language agent tasks. To address this, we present Reflection-Reinforced Self-Training (Re-ReST), which uses a reflector to refine low-quality generated samples during self-training. The reflector takes the agent’s output and feedback from an external environment (e.g., unit test results in code generation) to produce improved samples. This technique enhances the quality of inferior samples and efficiently enriches the self-training dataset with higher-quality samples. We conduct extensive experiments on open-source language agents across tasks, including multi-hop question answering, sequential decision-making, code generation, visual question answering, and text-to-image generation. The results demonstrate the effectiveness of self-training and Re-ReST in language agent tasks, with self-training improving baselines by 7.6% on HotpotQA and 28.4% on AlfWorld, and Re-ReST further boosting performance by 2.0% and 14.1%, respectively. Our studies also confirm the efficiency of using a reflector to generate high-quality samples for self-training. Moreover, we demonstrate a method to employ reflection during inference without ground-truth feedback, addressing the limitation of previous reflection work. Our code is released at https://github.com/PlusLabNLP/Re-ReST.
Re-ReST: Reflection-Reinforced Self-Training for Language Agents
Zi-Yi Dou, Cheng-Fu Yang , Xueqing Wu , Kai-Wei Chang , Nanyun Peng University of California, Los Angeles {zdou,cfyang,xueqing.wu,kwchang,violetpeng}@cs.ucla.edu
@@ 1 Introduction
Large language models (LLMs) (Kenton and Toutanova, 2019; Touvron et al., 2023; Achiam et al., 2023) have demonstrated potential in interacting with external environments and addressing practical interactive tasks, resulting in a new class — language agents (Nakano et al., 2021; Yao et al., 2022). Finetuning LLMs for agentic tasks has proven effective, yet existing works rely on data generated by stronger models (e.g., GPT-4) (Chen et al., 2023; Yin et al., 2024), which are not always available (e.g., to improve the strongest model).
Among the potential techniques to improve agents (Ouyang et al., 2022; Wang et al., 2023b; Li et al., 2024; Chen et al., 2024), self-training holds promise for enhancing agent performance for challenging agentic tasks. The self-training process typically involves refining the model by generating samples, assessing their quality through rewards, and updating the model by training on high-quality samples. Compared with existing agent training methods Chen et al. (2023); Yin et al. (2024), self-training can autonomously improve agents and reduce the discrepancy between the agent’s training data and its original predictions. Additionally, as in Figure 1, self-training can potentially allow for the development of performant agents within a fully open-source framework, without relying on closed-source, proprietary models. Given these benefits, we propose to investigate the use of self-training in language agents in this paper.
@@
However, one significant challenge for applying self-training in language agent tasks lies in the acquisition of high-quality samples to achieve good performance. Specifically, self-training requires a substantial amount of high-quality samples, while relying solely on model-generated samples can be inefficient, particularly for language agent tasks that demand multi-step reasoning and long-horizon planning. As a result, it is challenging to obtain good samples solely through sampling. Moreover, the common practice of discarding low-quality samples neglects their potential for improvement and effective utilization, thus limiting the overall efficacy of self-training methods.
To address these issues, we propose Reflection-Reinforced Self-Training (Re-ReST), which enhances the self-training algorithm using a reflection model. Re-ReST incorporates a reflector during self-training, which improves sample quality by utilizing environmental feedback such as execution successes and unit test outcomes. Specifically, the reflector transforms lower-quality samples into higher-quality ones, leveraging the capability of LLMs to self-improve when provided with accurate ground-truth feedback Huang et al. (2024). Consequently, it enriches the training dataset, enabling more effective bootstrapping. After training, only the agent model is used for inference, ensuring no additional computational burden during testing. Unlike existing self-reflection methods (Madaan et al., 2023; Shinn et al., 2023; Pan et al., 2023), Re-ReST only requires access to feedback during training, not during inference, making our setting more realistic and practical.
We conduct extensive experiments with open-source LLMs across a wide range of tasks, including multi-hop question answering, sequential decision-making, code generation, visual question answering, and text-to-image generation. Our results first demonstrate the potential of self-training in language agent tasks, showing improvements over few-shot baselines in long-horizon planning tasks, with gains of 7.6% on HotpotQA and 28.4% on AlfWorld. By incorporating Re-ReST, we further enhance performance significantly by 2.0% and 14.1% on HotpotQA and AlfWorld, respectively, achieving results better or comparable to models relying on commercial APIs. Ablation studies confirm the efficiency of the reflection model in generating high-quality self-training samples. Furthermore, we explore using our reflection model during inference with self-consistency decoding, which improves the model performance while alleviating the need for ground-truth feedback required by previous work Huang et al. (2024). Additionally, we demonstrate the application of our method in preference optimization objectives.
@@ 2 Method: Re-ReST
Self-Training.
Formally, given a dataset , self-training begins by using a base model to generate a pseudo-label for each instance . Subsequently, a subset of is selected based on a scoring function, and is finetuned on this selected subset. For language agents, we define the label as a trajectory comprising interleaved thoughts and actions, as described in ReAct Yao et al. (2022). We propose adopting the self-training paradigm by training language agents with their self-generated thought-action trajectories.
@@
Obtaining high-quality samples through self-sampling can be challenging, particularly for complex language agent tasks. To address this issue, we introduce Re-ReST, which aims to enhance the pseudo-label generation process in self-training for language agents. As illustrated in Figure 2, we propose improving low-quality samples using a reflection model with external feedback. We then enrich the self-training data by incorporating these corrected generations. This process generates high-quality samples efficiently by correcting low-quality ones with ground-truth feedback during training.
@@ 2.1 Components
The language agent is built upon a large language model (LLM) that is trained or prompted to generate thoughts and actions given a task. Formally, given an instance , the agent generates its output containing its actions. The agent can first generate its reasoning traces before outputting its actions, which has been demonstrated to improve the model performance and interpretability (Yao et al., 2022).
@@
The reflection model is also instantiated as an LLM, the goal of which is to improve the language agent’s generations given external feedback. We assume that during training, an external environment can evaluate a generated sample and provide feedback to the agent. The feedback can be a binary success status and/or error information. For example, in code generation tasks, the environment can execute the model-generated code on unit tests, providing information on whether the code has syntax errors and whether it can pass the unit tests. Having access to such an environment is important in our setting, as it has been shown that an LLM cannot perform self-correction without high-quality external feedback (Huang et al., 2024). The reflection model generates a corrected sample given the task information , the agent generation , and the environmental feedback . It can optionally first state its reasoning process (e.g., which specific actions could be corrected) before generating the corrected answer.) The use of the reflection model can improve self-training by finding good solutions efficiently because of the additional information provided (i.e., the agent’s previous trial and the environmental feedback.) We do not share the model parameters between the agent and reflector in this paper.
@@ 2.2 Data Generation
We then describe how we generate self-training data for the language agent . The data generation process involves two steps, including the initial generation step with the language agent itself and the reflection step with the reflector, and we obtain the agent-generated dataset and reflector-generated dataset from the two steps.
@@
As in the standard setup, given an instance , we sample generations from the current language agent model . Then, the environment scores the generation and provides feedback . If the score exceeds a threshold, we add the instance to to the training data . In practice, we observe that setting achieves a good balance between efficiency and effectiveness.
@@
The initial generation step only relies on the agent model itself to generate data. For a sampled generation , if the score does not pass the threshold, we will feed it to the reflection model for refinement. The reflector takes as inputs the task information , the agent’s prior generation , and the environmental feedback , and then generates the corrected sample The corrected sample will also be evaluated by the environment and we will add it to the reflector-generated training dataset if its score exceeds the threshold. While the reflection procedure can be iteratively applied multiple times as per Shinn et al. (2023), in this study, we limit this process to a single iteration for the sake of efficiency. This means that each generated sample is allowed a maximum of one refined counterpart .
@@ 2.3 Model Training and Inference
We first train the reflector parameterized by and then use the trained reflector to generate the reflection data . Afterward, we combine and the agent’s self-generated data to train the agent model parameterized by .
@@
While base LLMs can perform self-reflection or self-correction without any finetuning given ground-truth feedback (Shinn et al., 2023), we propose to further improve its reflection ability with the self-generated data. First, from the initial generation step, we obtain multiple generations from the agent model . For each correct generation and incorrect generation with its environmental feedback in , we will add the instance to the agent-generated dataset for reflector training. In addition, the reflector generates its self-training dataset in a zero-shot manner similar to the agent initial generation step. Combining the two generated datasets, we train the reflector on with the standard maximum log-likelihood objective first before generating the training data for the language agent:
@@
After we have the base language agent to generate the self-training data and the improved reflector to generate the reflector-generated data , we train the language agent jointly on :
Besides the maximum log-likelihood objective, because the reflection training and data generation process involves the use of preference pairs, it is natural to use preference optimization objectives such as DPO (Rafailov et al., 2023) for training, which we will discuss in the experiment section.
@@
During inference, accessing high-quality environmental feedback is often challenging, which can cause inference-time self-reflection algorithms to fail (Huang et al., 2024). Therefore, we only have the agent directly output generations without the reflector during inference. This approach eliminates the need for feedback and avoids any additional computational overhead. A potential method to integrate the reflector into the inference process involves first training a scorer to evaluate the agent’s output. If the score falls below a certain threshold, self-correction can then be performed, which we leave as a future direction. Additionally, we propose performing reflection regardless of environmental feedback and employing self-consistency to derive the final results from both the agent’s outputs and the reflector’s outputs, as shown in the experiment section.
@@ 3 Experiments
We experiment with multi-hop reasoning, sequential decision-making, code generation, visual question answering, and text-to-image generation. We present the experimental settings and results for each task. In all our experiments, we advocate for the use of open-source models and aim to avoid black-box, closed-source commercial models whenever possible.
@@ 3.1 Multi-Hop Reasoning
We use the HotpotQA dataset (Yang et al., 2018), a well-established question-answering dataset featuring multi-hop reasoning and knowledge retrieval. It is constructed based on Wikipedia and an agent needs to retrieve and reason over multiple supporting documents to answer a question. We sample 5,000 training instances randomly for self-training and 500 instances from the development set for evaluation as in Chen et al. (2023).
@@
We use 2-shot prompting for few-shot agent and reflector data generation as in Shinn et al. (2023). For each training instance, the agent model samples 3 generations. The generation is evaluated with the exact match metric (i.e., if the generated answer is exactly the same as the ground-truth answer). The retrieval and evaluation results are given to the reflector as the environmental feedback for self-correction. We use Low-Rank Adaptation (LoRA) (Hu et al., 2022) for training the language models for efficiency. The agent and reflector models are trained for 3 epochs with a learning rate of 3e-4.
| Model | QA Tool | #Train Data | EM |
@@
| Llama-2-13B ReAct-Based Agents | |||
| Few-Shot | WikipediaAPI | - | 20.0 |
@@
| Llama-2-13B ReAct-Based Agents w/ GPT-4-Generated Data | |||
| FireAct (Chen et al., 2023) | SerpAPI | 0/0.5k | 34.4 |
| LUMOS (Yin et al., 2024) | GPT-3.5 | 0/20k | 31.4 |
| LUMOS (Yin et al., 2024) | GPT-4 | 0/20k | 36.3 |
| FireAct | WikipediaAPI | 0/0.5k | 32.2 |
| Self-Training | WikipediaAPI | 2.5k/0.5k | 34.2 |
| Re-ReST | WikipediaAPI | 3k/0.5k | 35.8 |
| Llama-3-8B ReAct-Based Agents | |||
| Few-Shot | WikipediaAPI | - | 30.0 |
@@
We list the main results in Table 1. As shown in the table, self-training can significantly improve the model performance from an EM score of 20.0% to 27.6% for Llama-2 and from 30.0% to 34.4% for Llama-3. However, only 37.1% and 48.3% of the training instances are correctly solved by the agent model and are used for self-training respectively. By integrating our reflector model into the process, the agent can solve more training instances and thus have more data for training the agent model, increasing the EM scores significantly. In addition to our implemented models, following previous work (FireAct (Chen et al., 2023) and LUMOS Yin et al. (2024)) that use GPT-3.5/4 for data generation and model finetuning, we employ GPT-4 to generate 0.5k instances and first train the agents with the GPT-4 generated data before self-training. Results demonstrate that 1) self-training is a stronger baseline than FireAct under a fair setting where the same QA tool is used; 2) we can achieve comparable or better performance of our model than these methods, even though both of them use strong knowledge retrieval models (i.e., SerpAPI222https://serpapi.com/ for FireAct and GPT-4 for LUMOS), which are costly and non-scalable. By contrast, we use the free Wikipedia API.